OpenStudy (littlediva12459):

Determine which pair of points has a positive slope. (5, –4), (–2, 1) (6, –10), (2, 10) (–10, –2), (6, 6) (5, –1), (–6, 6)

OpenStudy (vladthepro):

First of all, you have to use the equation (Y2-Y1) / (X2-X1)

OpenStudy (littlediva12459):

Would it be (–10, –2), (6, 6)

OpenStudy (vladthepro):

\[\frac{ (-4-1) }{ (5+2) } = \frac{ -3 }{ 7 }\]

OpenStudy (vladthepro):

Yes, that is correct :)

OpenStudy (littlediva12459):

Thank you

OpenStudy (vladthepro):

Anytime ^.^